Two small planes start from the same point and fly in opposite directions. The first plane is flying 70 mph slower than the second plane. In 3 h, the planes are 870 mi apart. Find the rate of each plane.
slower plane:
faster plane:

Answer by CubeyThePenguin(3113)

speed of fast plane = x mph
speed of slow plane = (x-70) mph

distance between planes after h hours: (2x-70)h miles

h = 3 ----> 3(2x - 70) = 870
2x - 70 = 290
2x = 360
x = 180

slower plane = 110 mph
faster plane = 180 mph